# What are the Essential Factors in Crafting Effective Long Context Multi-Hop Instruction Datasets? Insights and Best Practices

### v1

@@ Abstract
Recent advancements in large language models (LLMs) with extended context windows have significantly improved tasks such as information extraction, question answering, and complex planning scenarios. In order to achieve success in long context tasks, a large amount of work has been done to enhance the long context capabilities of the model through synthetic data. Existing methods typically utilize the Self-Instruct framework to generate instruction tuning data for better long context capability improvement. However, our preliminary experiments indicate that less than 35% of generated samples are multi-hop, and more than 40% exhibit poor quality, limiting comprehensive understanding and further research. To improve the quality of synthetic data, we propose the Multi-agent Interactive Multi-hop Generation (MIMG) framework, incorporating a Quality Verification Agent, a Single-hop Question Generation Agent, a Multiple Question Sampling Strategy, and a Multi-hop Question Merger Agent. This framework improves the data quality, with the proportion of high-quality, multi-hop, and diverse data exceeding 85%. Furthermore, we systematically investigate strategies for document selection, question merging, and validation techniques through extensive experiments across various models. Our findings show that our synthetic high-quality long-context instruction data significantly enhances model performance, even surpassing models trained on larger amounts of human-annotated data. Our code is available at: https://github.com/WowCZ/LongMIT.
@@ 1 Introduction
Recently, large language models (LLMs) with long context windows have significantly improved tasks such as information extraction, question answering, and even complex planning scenarios (Liu et al., 2024a; Bai et al., 2024b; Hu et al., 2023; 2024; Xu et al., 2024b). Research on developing long-context LLMs has predominantly focused on extending the context window (Ding et al., 2024; Jin et al., 2024; Peng et al., 2024). Nevertheless, in practical applications, simply expanding the context window proves inadequate (Hsieh et al., 2024; Huang, 2024). There is a pressing need for training to optimize utilization of long context (Zhang et al., 2024), especially in instruction tuning (Fu et al., 2024b).
In the instruction tuning phase, a large amount of high-quality long context instruction data is required. However, high-quality long-context instruction data is extremely difficult to obtain, and the annotation cost is extremely high compared to short-context data (Bai et al., 2024b; Xiong et al., 2024). Therefore, Xiong et al. (2023) and Bai et al. (2024a) have begun to explore the use of LLM to generate instruction tuning data through the Self-Instruct framework (Wang et al., 2023b), which addresses the issue of long-context instruction tuning data scarcity.
@@
Figure 1: Comparison between traditional self-instruct-based data synthesis method and our Multi-agent Interactive Multi-hop Generation (MIMG) data synthesis framework.
Despite addressing data scarcity, existing works still struggle to synthesize high-quality multi-hop instruction tuning data. This gap arises from inadequate focus on the synthesis process and the factors that affect data effectiveness. As shown in Figure 1, our preliminary manual annotation experiments reveal that direct self-instruction generates less than 35% of multi-hop samples, with high-quality samples accounting for only around 60%. Moreover, sample diversity is concerning, with more than 45% of samples with semantic duplication and lack of diversity. These limitations hinder a comprehensive understanding of the mechanisms and impede further research in this field.
Motivated by these challenges, this paper systematically investigates the research question: What are the essential factors in crafting effective long-context multi-hop instruction datasets? To address this, we propose a Multi-agent Interactive Multi-hop Generation (MIMG) framework. First, to enhance the quality of long-context instruction data, we introduce a novel Quality Verification Agent, which prompts LLMs to automatically verify the quality of generated samples. Second, to incorporate multi-hop reasoning, we develop a Single-hop Question Generation Agent, followed by a Multi-hop Question Merging Agent for stepwise synthesis of multi-hop problems. Finally, to ensure diversity, we implement multiple question sampling strategies within the Single-hop Question Generation Agent to minimize repetition and promote variety in the multi-hop questions. As shown in Figure 1, our method significantly improves data quality, with over 85% of the data being multi-hop, high-quality, and non-duplicative.
To understand the best practices for crafting long-context instruction data, we systematically investigate several sub-questions, including strategies for document selection, the impact of different question merging approaches, and the effectiveness of various validation techniques. We conduct extensive experiments using 17 strategies across 10 domains and 5 models. Our findings show that the proposed methods significantly enhance data quality and improve the performance of long-context understanding models. Moreover, models trained on our synthetic high-quality long-text instruction data can achieve at least 7.54% on average improvement, which even surpasses the performance of models trained on a larger volume of human-annotated data.
The main contributions of our work are as follows:
  • An Extensive Exploration of Best Practices: Our study examines a series of strategies for generating high-quality multi-hop instruction data, identifying key factors and best practices that impact long-context data quality.
  • A Novel Data Synthesis Framework: We introduce a Multi-Agent Interactive Multi-hop Generation (MIMG) framework to synthesize multi-hop instruction data, enhancing data quality.
  • A Large Long-Context Instruction Dataset for Effectively Enhanced Long-Context Utilization: Our synthetic dataset (LongMIT) proves effective across numerous models, outperforming even larger human-labeled instruction datasets.
@@
Figure 2: The overall process of our Multi-agent Interactive Multi-hop Generation (MIMG) data synthesis framework.
@@ 2 Framework
Our proposed framework consists of four main components: Quality Verification Agent (§$\S$ 2.1), Single-hop Question Generation Agent (§$\S$ 2.2), Multiple Question Sampling (§$\S$ 2.3), and Multi-hop Question Merging Agent (§$\S$ 2.4). Specifically, each component is meticulously designed to contribute to the generation of high-quality multi-hop questions and answers from given documents. First, the Quality Verification Agent ensures the data quality generated by other agents. The Single-hop Question Generation Agent then creates initial single-hop questions, focusing on simple, direct inquiries. Next, Multiple Question Sampling strategies expand on this by sampling questions that cover various aspects of the document, enhancing multi-hop instruction generation. Finally, the Multi-hop Question Merging Agent integrates these single-hop questions into coherent multi-hop questions, requiring information synthesis from multiple document parts, ensuring deeper understanding and engagement. The detailed architecture is illustrated in Figure 2.
@@ 2.1 Quality Verification Agent
The first module in our framework is Quality Verification Agent, which ensures that the generated questions and answers meet a certain standard of quality. This component involves two main processes:
@@
This includes additional heuristic strategies to judge which samples should be contained as high-quality data. Specifically, we utilize two wide-used verification strategies:
  • Scoring: We prompt LLMs to generate continuous scores, manually set a more reliable threshold score based on the validation set, and set those exceeding the threshold score as high-quality data. Formally, given a sample $s$, we select the high-quality data as follows:
    $\mathcal{V}(s|\mathcal{M})=\begin{cases}\texttt{Approved}&\texttt{Score}(s|\mathcal{M})>\theta;\\ \texttt{Rejected}&\texttt{Score}(s|\mathcal{M})\leq\theta,\end{cases}$ (1)
    where $\texttt{Score}(s|\mathcal{M})$ represents the model score of sample $s$ based on model $\mathcal{M}$, and $\theta$ is the threshold.
  • Classification: We prompt LLMs to generate binary classification and select those classified as high-quality data. Formally, given a sample $s$, we select the high-quality data as follows:
    $\mathcal{V}(s|\mathcal{M})=\begin{cases}\texttt{Approved}&\texttt{Class}(s|\mathcal{M})=1;\\ \texttt{Rejected}&\texttt{Class}(s|\mathcal{M})=0,\end{cases}$ (2)
    where $\texttt{Class}(s|\mathcal{M})$ represents the binary classification process of sample $s$.
@@
  • Auxiliary Generation Information: We enable the model to provide more reasoning rationale during output generation and observe whether this improves the robustness and accuracy of the verification process.
@@ 2.2 Single-hop Question Generation Agent
This phase generates single-hop questions and answers from individual documents, encompassing the following components:
@@
This component utilizes a robust LLM to generate valid and relevant single-hop questions and answers from each document. Multiple questions and answers are produced per document to ensure a diverse foundation for multi-hop question development. We thoroughly examine various LLMs, including both open-source and close-source models, across different scales.
@@
The strategy employs a structured approach to extract potential questions from the text, using the following techniques:
  • Rationale-based Question Generation: Chain-of-Thought (CoT) prompting (Wei et al., 2022) has been recognized for its role in improving performance on long-text tasks (Li et al., 2024). Building on this, our study investigates whether generating questions from a long document, supported by rationale, can enhance the understanding of the document’s inherent reasoning.
  • Question-Answering Generation Order: Furthermore, we aim to evaluate whether the sequence of generating questions and answers impacts the overall effectiveness. Specifically, generating the question prior to the answer may reduce the reasoning complexity and improve the quality of the model’s output compared to a simultaneous generation approach.
@@ 2.3 Multiple Question Sampling
In order to further optimize the diversity of generated samples, we introduce Multiple Question Sampling strategy to create multi-hop questions by sampling and combining questions from multiple questions and documents. It mainly involves the following two strategies:
@@
This strategy identifies relevant questions and documents for multi-hop question creation. Using relevance sampling, a question semantic relevance matrix is generated, assessing the semantic connections between questions across different documents and guiding the sampling process. The strategy includes:
  • Probability-Based Sampling: This method evaluates document relevance based on the probability and occurrence of specific keywords related to the questions, like BM25 (Robertson et al., 1995; 2009), and LDA (Hoffman et al., 2010).
@@
Based on the relevance matrix, the most related questions are selected for merging. This involves choosing questions that are both relevant and complementary, ensuring that the resulting multi-hop questions are coherent and contextually rich. The strategy includes:
  • Intra-Document Sampling: This strategy focuses on selecting questions within the same document to ensure internal coherent multi-hop data.
@@ 2.4 Multi-hop Question Merging Agent
We utilize LLM to combine the sampled questions and answers into meaningful multi-hop questions and answers. The model leverages context and semantic understanding to ensure that the merged questions are logically consistent and contextually accurate. The backbone includes 5 classic LLMs.
@@
This includes rules and heuristics to ensure the merged questions are logically consistent and contextually accurate. The strategy includes:
  • Document-Based Merging: To further reduce input tokens, we explore whether long documents need to be added to large model inputs to enhance merging performance. Formally, the merging process can be represented as:
    ${Q_{m}}=\mathcal{M}(Q_{1},Q_{2},\ldots,Q_{n}|C).$ (6)
    where $Q_{1},Q_{2},\ldots,Q_{n}$ are the sampled single-hop questions, and $Q_{m}$ represents the merged multi-hop question. $C$ denotes context whether utilize document.
  • Rationale-Based Merging: This method leverages the underlying rationale or reasoning behind the original questions to guide their integration, ensuring that the combined question preserves the intended meaning and context of the individual components. Formally, this merging process can be expressed as:
    $R\oplus{Q_{m}}=\mathcal{M}(Q_{1},Q_{2},\ldots,Q_{n}).$ (7)
    where $R$ represents the rationale or underlying reasoning, and $\oplus$ denotes the connector vocabulary in generated response.
Furthermore, we explore the creation of both intra-document and inter-document multi-hop instruction samples for different scenarios.
@@ 3.1.1 Verification Strategy
Currently, the most widely employed strategies for model verification are scoring and direct classification. We evaluated the consistency and precision of both approaches by comparing them with human annotations in the sample analysis of data generated from long contexts.
@@
As shown in Figure 3 (a), the scoring strategy shows significantly higher kappa and precision scores compared to binary quality classification. This statistical improvement suggests that scoring better captures the nuances of human judgments. This observation aligns with findings in short-context scenarios (Fu et al., 2024a), reinforcing the generalizability of scoring strategies across different lengths of textual data.
@@
As depicted in Figure 3 (a), in contrast to their performance in short-context verification (Wang et al., 2023a; Fu et al., 2024a), LLMs demonstrate minimal agreement with human annotators in long-context scenarios, reflected in low kappa scores. This suggests challenges in maintaining annotation consistency due to the cognitive load and interpretative variations over extensive information. Despite this, as demonstrated in Figure 3 (b), LLMs consistently achieve nearly perfect precision, indicating robust capability in identifying and selecting relevant data. This distinction underscores the potential of LLMs as effective tools for data filtering and prioritization in long-context environments, contrasting their role as accurate annotators in short-context scenarios.
Figure 4: Performance of different models for generating single-hop questions.
Scoring can alleviate the long context bias but classification not.
Additionally, we examine why classification strategies underperform in long contexts by assessing verifier precision across varying context lengths. As depicted in Figure 4, the scoring strategy demonstrates superior precision and greater robustness in extended contexts, explaining the poor performance of classification in these verification scenarios.
Based on the conclusions drawn from our prior analyses, the subsequent experiments will be conducted within the framework of the Scoring strategy. The precision of the verifier will serve as the metric for evaluating its quality, while the quality of the data will be assessed by examining the data retention ratio.

### v2

@@ Abstract
Recent advancements in large language models (LLMs) with extended context windows have significantly improved various tasks. To improve long-context capabilities, much work focuses on augmenting LLM’s capabilities with synthetic data. Existing methods often leverage the Self-Instruct framework to generate long-context instruction-tuning data. However, our preliminary experiments show that fewer than 35% of samples generated by Qwen-272B are multi-hop, and over 40% exhibit poor quality, limiting comprehensive understanding and further research. To address this, we propose the Multi-agent Interactive Multi-hop Generation (MIMG) framework, which integrates a quality verification agent, a single-hop question generation agent, a multiple question sampling strategy, and a multi-hop question merger agent. This framework significantly improves data quality, with high-quality, multi-hop, and diverse data. Furthermore, we conduct a thorough analysis of document selection, question merging, and validation techniques through extensive experiments across various models. Our results demonstrate that synthetic high-quality long-context instruction data can enhance model performance, surpassing even models trained on larger amounts of human-annotated data. Our code and relevant data are available at: https://github.com/WowCZ/LongMIT.
What are the Essential Factors in Crafting Effective Long Context Multi-Hop Instruction Datasets? Insights and Best Practices
Zhi Chenthanks:   Equal Contribution   Qiguang Chen11footnotemark: 1    Libo Qinthanks:   Corresponding Author    Qipeng Guo    Haijun Lv
Yicheng Zou   Hang Yan    Kai Chen    Dahua Lin Shanghai Artificial Intelligence Laboratory Research Center for Social Computing and Interactive Robotics Harbin Institute of Technology School of Computer Science and Engineering, Central South University chenzhi@pjlab.org.cn, qgchen@ir.hit.edu.cn
@@ 1 Introduction
Figure 1: Comparison between traditional self-instruct-based data synthesis method and our Multi-agent Interactive Multi-hop Generation (MIMG) framework, where all data are generated by Qwen-272B (Yang et al., 2024).
@@
Figure 2: The overall process of Multi-agent Interactive Multi-hop Generation (MIMG) data synthesis framework.
Recently, large language models (LLMs) with long-context windows have significantly improved tasks such as information extraction, question answering, and even complex planning scenarios (Liu et al., 2024a; Bai et al., 2024b; Hu et al., 2025; Xu et al., 2024b). Research on developing long-context LLMs has predominantly focused on extending the context window (Ding et al., 2024; Jin et al., 2024; Peng et al., 2024). Nevertheless, in practical applications, merely expanding the context window is insufficient for effectively utilizing long-context (Hsieh et al., 2024; Huang, 2024), which presses a need for training to optimize utilization of long-context (Zhang et al., 2024a), especially in instruction-tuning (IT) (Fu et al., 2024b). In the IT phase, a large amount of high-quality long-context IT data is required. However, acquiring such data is challenging, with annotation costs significantly higher than those for short-context data (Bai et al., 2024b; Xiong et al., 2024). To address this, Xiong et al. (2023) and Bai et al. (2024a) have explored leveraging LLMs to generate IT data using the Self-Instruct framework (Wang et al., 2023b), thereby mitigating the scarcity of long-context IT data.
Moreover, the challenge often lies not in extracting single-hop information, but in integrating multiple hops of data from long contexts to derive complex conclusions. However, existing studies struggle to generate high-quality, multi-hop IT data, primarily due to insufficient focus on the data synthesis process and the factors influencing data effectiveness. As illustrated in Figure 1 (a), our preliminary manual annotation experiments show that direct self-instruction yields less than 35% multi-hop samples, with high-quality examples representing only 60%. Additionally, sample diversity remains problematic, with over 45% of the samples exhibiting semantic duplication. These issues hinder comprehensive understanding and further advancement in this domain [^1] See Appendix A for all detailed metric description..
Motivated by these challenges, this paper systematically investigates the research question: What are the essential factors in crafting effective long-context multi-hop instruction datasets? To address this, inspired by recent advancement on Agent (Hu et al., 2023; Wang et al., 2024), we propose a Multi-agent Interactive Multi-hop Generation (MIMG) framework. First, to ensure data quality, a Quality Verification Agent is introduced to evaluate the quality of long-context samples throughout the process. Second, for multi-hop reasoning, a Single-hop Question Generation Agent will be followed by a Multi-hop Question Merging Agent for stepwise synthesis of multi-hop queries. Finally, to ensure diversity, Multiple Question Sampling strategies are proposed to reduce redundancy and promote variety. To comprehensively examine the factors in long-context multi-hop data creation, we conduct extensive experiments, applying 17 strategies across 10 domains and 5 LLMs. As shown in Figure 1 (b), our method significantly improves data quality, yielding over 85% multi-hop, high-quality, and non-duplicative samples. Notably, LLMs trained on the synthetic high-quality data show an average improvement of 7.54%, even surpassing those LLMs trained on larger human-annotated datasets.
Overall, the main contributions are as follows:
  • We systematically explore strategies for generating high-quality multi-hop instruction data to identify unexplored but critical factors, that influence the quality of long-context data. These factors include scoring verifiers, question-then-answer generation, question-based sampling, and question-answer merging strategies.
  • We introduce the Multi-Agent Interactive Multi-hop Generation (MIMG) framework, which enhances the quality and relevance of synthesized data through multiple agent interactions.
  • Our synthetic dataset, LongMIT, has shown superior performance across various long-context datasets. It not only improves long-context utilization but also surpasses larger human-labeled datasets, demonstrating its practical impact on advancing long-context LLMs.
@@ 2 Framework
Our framework consists of 4 main components: quality verification agent (QVA; §$\S$ 2.1), single-hop question generation agent (SQGA; §$\S$ 2.2), multiple question sampling (MQS; §$\S$ 2.3), and multi-hop question merging agent (MQMA; §$\S$ 2.4). Specifically, the QVA is first designed as a validator to control and supervise the data quality at each stage. The SQGA then generates simple and direct single-hop questions. Next, MQS strategies expand on this by sampling questions that cover various documents, enhancing multi-hop instruction generation. Finally, the MQMA integrates these single-hop questions into the multi-hop questions, requiring multiple document information. The detailed architecture is illustrated in Figure 2.
@@ 2.1 Quality Verification Agent
The first module is the Quality Verification Agent, which globally supervises and ensures the quality of generated samples. This component involves two main processes:
@@
This includes additional heuristic strategies to judge which samples should be contained as high-quality data. Specifically, we explore two widely-used verification strategies:
  • Scoring: We prompt LLMs to assign continuous scores and determine a threshold using the validation set to filter high-quality data. Formally, given a sample $s$, the selection criterion is:
    $\!\!\mathcal{V}(s|\mathcal{M})\!=\!\begin{cases}\texttt{Approved}&\mathcal{F}_{S}(s|\mathcal{M})\!>\!\theta;\\ \texttt{Rejected}&\mathcal{F}_{S}(s|\mathcal{M})\!\leq\!\theta,\end{cases}$ (1)
    where $\mathcal{F}_{S}(s|\mathcal{M})$ represents the score of sample $s$ based on model $\mathcal{M}$, and $\theta$ is the threshold.
  • Classification: We prompt LLMs to perform binary classification and retain only samples classified as high-quality. Formally, given a sample $s$, the selection criterion is:
    $\!\!\mathcal{V}(s|\mathcal{M})\!=\!\begin{cases}\texttt{Approved}&\mathcal{F}_{C}(s|\mathcal{M})\!=\!1;\\ \texttt{Rejected}&\mathcal{F}_{C}(s|\mathcal{M})\!=\!0,\end{cases}$ (2)
    where $\mathcal{F}_{C}(s|\mathcal{M})$ represents the binary classification process of sample $s$.
@@
  • Auxiliary Generation Information: We enable the model to provide reasoning rationale during output generation and observe its effectiveness.
@@ 2.2 Single-hop Question Generation Agent
This phase generates single-hop questions and answers from individual documents through the following components:
@@
We need a robust LLM to generate multiple single-hop questions and answers per document, ensuring diversity for data synthesis. Therefore, we evaluate various LLMs, both open- and closed-source, across different scales.
@@
Our strategy employs structured techniques to extract potential questions:
  • Rationale-based Question Generation: Chain-of-Thought (CoT) can effectively enhances performance on long-context tasks (Li et al., 2024). Hence, we explore whether generating rationale-supported questions improves the model’s grasp of a document’s reasoning structure.
  • Question-Answering Generation Order: Furthermore, we assess whether order affects output quality. Generating questions before answers may simplify reasoning and enhance output clarity compared to simultaneous generation.
@@ 2.3 Multiple Question Sampling
To further optimize the diversity of generated samples, we introduce MQS strategy, which constructs multi-hop questions by sampling and combining questions from various documents. This approach consists of two key strategies:
@@
This strategy selects relevant questions and documents for multi-hop question generation. By leveraging relevance sampling, it constructs a question-semantic relevance matrix to assess semantic connections across different documents, guiding the sampling process. The strategy comprises:
  • Probability-Based Sampling: This method samples data based on probability-based document relevance, which is calculated as the frequency of keywords related to the questions, like BM25 (Robertson et al., 1995, 2009), and LDA (Hoffman et al., 2010).
@@
Based on the relevance matrix, the most related questions should be selected to form a coherent, contextually rich multi-hop question. The strategy includes:
  • Intra-Document Sampling: It focuses on selecting questions within the same document to ensure internal coherent multi-hop data.
@@ 2.4 Multi-hop Question Merging Agent
We should employ LLMs to synthesize sampled questions and answers into meaningful multi-hop queries. To investigate this, we conduct the following exploration:
@@
This strategy applies rules and heuristics to maintain logical and contextual consistency. It includes:
  • Document-Based Merging: To further reduce input tokens, we explore whether incorporating long documents enhances merging performance. Formally, the process is:
    ${Q_{m}}=\mathcal{M}(Q_{1},Q_{2},\ldots,Q_{n}|C),$ (6)
    where $Q_{1},Q_{2},\ldots,Q_{n}$ are sampled single-hop questions, $Q_{m}$ is the merged multi-hop query, and $C$ denotes whether documents are used.
  • Rationale-Based Merging: To preserve the meaning and context of the individual questions, this method leverages the reasoning rationale behind the original questions to guide their merging process, which can be expressed as:
    $R\oplus{Q_{m}}=\mathcal{M}(Q_{1},Q_{2},\ldots,Q_{n}),$ (7)
    where $R$ represents the underlying rationale, and $\oplus$ denotes the connecting elements in the generated response.
Additionally, we explore intra-document and inter-document multi-hop instruction samples for diverse scenarios.
@@ 3 Exploration
This section examines the framework components aimed at enhancing data quality, including verification strategies in QVA (§3.1), generation strategies in SQGA (§3.2), sampling approaches in MQS (§3.3), and merging strategies in MQMA (§3.4).
@@ 3.1.1 Verification Strategy
Currently, the most commonly used model verification strategies are scoring and direct classification. We assessed the consistency and accuracy of both methods by comparing them with human annotations in a sample analysis of long-context data.
@@
As shown in Figure 3 (a), the scoring method significantly outperforms binary classification, yielding higher kappa and precision scores. This statistical improvement indicates that scoring better captures the subtleties of human judgment. These findings align with research in short-context scenarios (Fu et al., 2024a), highlighting the broader applicability of scoring methods across different context lengths.
@@
As depicted in Figure 3 (a), unlike their performance in short-context verification (Wang et al., 2023a; Fu et al., 2024a), LLMs show minimal agreement with human annotators in long-context situations, reflected in low kappa scores. This suggests challenges in maintaining consistent annotations due to cognitive load and interpretive variations over extended data. However, as Figure 3 (b) shows, LLMs maintain near-perfect precision, demonstrating their strong ability to identify and select relevant information. This highlights LLMs’ potential as effective tools for data filtering and prioritization in long-context environments, in contrast to their role as accurate annotators in short-context settings.
Scoring alleviates the long context bias but classification does not.
We further examine why classification performs poorly in long-context scenarios by analyzing precision across different context lengths. Figure 3 (c) illustrates that scoring consistently achieves higher precision and robustness in longer contexts, explaining the suboptimal performance of classification in these cases. Based on these findings, subsequent experiments will adopt the scoring strategy, using verifier precision and the data retention ratio to assess data quality. More discussion can be seen in Appendix B.